Grade 12th pass9 grade science

Express 15 m/s in km/h.

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Yash Patari
4 Years ago
dear student
We know that 1km = 1000m
1 hr = 60 minutes = 3600sec
so km/hr = 1000/3600 = 5/18

Hence we multiply by 5/18 to convert km/hr to m/s

Here we have to convert m/s to km/hr
so we will do the reverse that is 18/5

15m/s = 15 × 18/5 = 3 × 18 = 54


Hence your answer is 54km/hr
 
thanks